Table 1: Car Ownership Satisfaction
Result (Country) Malaysia (Country) USA Country (China)
Satisfied 700 1200 1700
Not Satisfied 1300 800 300

a) Based on Table 1, identify the type of variables of the Country and Result.

b) Construct a relative frequency distribution (in decimal) for each of the result.
1
Expert's answer
2019-05-15T13:42:00-0400

a) Country - Nominal.

Result - Nominal.

b)

Malaysia USA China

Satisfied 700/3600=0.1944 1200/3600=0.3333 1700/3699=0.4722

Not satisfied 1300/2400=0.5417 800/2400=0.3333 300/3600=0.1250